# LLM Trading: Analysis of LLM Agent Behavior in Experimental Asset Markets

### v1

@@ Abstract
This paper explores how Large Language Models (LLMs) behave in a classic experimental finance paradigm widely known for eliciting bubbles and crashes in human participants. We adapt an established trading design—where traders buy and sell a risky asset with a known fundamental value—and introduce several LLM-based agents, both in single-model markets (all traders are instances of the same LLM) and in mixed-model “battle royale” settings (multiple LLMs competing in the same market). Our findings reveal that LLMs generally exhibit a “textbook-rational” approach, pricing the asset near its fundamental value, and show only a muted tendency toward bubble formation. Further analyses indicate that LLM-based agents display less trading strategy variance in contrast to humans. Taken together, these results highlight the risk of relying on LLM-only data to replicate human-driven market phenomena, as key behavioral features—such as large, emergent bubbles—were not robustly reproduced. While LLMs clearly possess the capacity for strategic decision-making, their relative consistency and rationality suggest that they do not accurately mimic human market dynamics.
@@ 1 Introduction
Financial decision-making is a complex and multifaceted process shaped by the interplay of risk, capital, and time. Traditional frameworks emphasize rational trade-offs, but human behavior often deviates due to biases such as perceptual ambiguities, overconfidence and herding behavior, these biases can cause short-term buying pressure and the over-inflation of asset values, culminating in market bubbles. These patterns highlight the persistent challenge of achieving optimal outcomes in dynamic and uncertain financial environments.
While there is extensive literature on human behavior in experimental finance settings, little work has been done on how large language models (LLMs) perform in these paradigms. Do LLMs exhibit similar behavioral biases, are they more ”rational” than humans, or do they fall victim to different biases not seen in humans. As LLMs increasingly contribute to domains previously managed by humans it is becoming more important to understand the tendencies and biases that these models exhibit in financial settings.
Previous studies suggest that interactions with agents differ significantly from those with humans (Anlló et al. (2024)), while others propose that LLM behavior could supplement human studies (Leng (2024); Horton (2023)). This paper investigates both claims by exploring how LLMs conduct themselves in experimental finance markets and their similarity to human traders.
@@
LLM markets behave much more ”rationally” than humans; they generally trade much closer to fundamental value. LLM markets are not similar to human markets and should not be used to mimic or pilot human data. In a competitive setting, the spread between the highest and lowest performing models is fairly moderate—indicating that while some agents capitalize on slightly better trading opportunities, the market does not produce extreme winners or losers.
@@ 2 Background
LLMs in Strategic and Financial Environments LLMs have demonstrated emergent capabilities in replicating human-like behavior in game theory and strategic environments. Brookins & DeBacker (2024) explored GPT-3.5’s cooperative tendencies in the Prisoner’s Dilemma, finding that it often exceeds human preferences for fairness. Horton (2023) investigated LLMs in public goods games, observing their ability to balance altruistic and selfish strategies depending on contextual prompts. Other studies have delved into multi-agent strategic interactions. Akata et al. (2023) examined GPT-4’s ability to adaptively switch between competitive and cooperative strategies in iterated games, revealing its capacity to engage in long-term planning.
In finance, LLMs have been applied to both experimental and real-world scenarios. Narayan et al. (2023) explored LLMs’ ability to process financial sentiment and predict market movements. Fatouros et al. (2024) introduced MarketSenseAI, leveraging GPT-4 for stock selection illustrating LLMs’ utility in navigating complex financial data. While these studies demonstrate the promise of LLMs in financial settings, there is still a need to understand how these agents perform in experimental financial settings where strategic interactions drive outcomes.
Experimental Finance The study of market dynamics in controlled environments has revealed critical insights into behavioral biases and inefficiencies. For instance, Smith et al. (1988) demonstrated the formation of bubbles and crashes in experimental markets, even under conditions of perfect information. Subsequent work by Bostian & Holt (2009) highlighted how individual behaviors, such as overconfidence and herding, contribute to these phenomena. Incorporating AI agents into such paradigms provides a unique opportunity to evaluate whether these systems replicate, suppress, or amplify such market inefficiencies.
@@ 3 Experiment
We focus on a classic experimental finance paradigm with an easily computable fundamental value, the true mathematical price of an asset; see Equation  1. The experimental design involves an experimental asset market where subjects trade a risky asset in an open-call market for 30 periods (Smith et al. (2014); Holt et al. (2017)) operated using the oTree experiment platform (Chen et al. (2016)). Participants are given 4 stocks and 100 units of cash. Participants have two options: either invest cash in the risky asset, which pays out dividends of either 0.4 or 1 with equal probability or hold cash which generates 5% interest rate per trading period. More details of the experiment can be found in Appendix A.
At the conclusion of the experiment, the risky asset is redeemed for a fixed value of 14 units, representing the fundamental price, calculated by the formula:
$\frac{\mathbb{E}[\text{Dividend}]}{\mathbb{E}[\text{Interest Rate}]}=\frac{0.7}{0.05}=14$ (1)
Note our experiment was designed such that this fundamental value remains constant throughout the entirety of the experiment, enabling us to objectively classify when the market is in a ”bubble” state Bostian et al. (2005).
In adapting the task to LLM-based agents, our primary goal was to best emulate the experiment presented to human subjects. The description of the task setting mirrors the original instruction video, shown to human participants. In every round of the experiment, each agent simultaneously submits orders to the market as well as expectations regarding current and future market prices.
A common challenge in adapting multi-round games to LLM-based players is the persistence of memory between rounds, which is critical to learning and adapting strategies. Building off previous works (Fish et al. (2024), we allow agents to communicate with future versions of themselves by allowing them to read and write their own ”Insights” and Thoughts” at each round. This structure incorporates Chain-of-Thought (CoT) reasoning (Wei et al. (2022)), which has been shown to be helpful in eliciting the reasoning ability of LLMs, and allows for an interpretable window into how agents perceive and wish to interact with the market. Further experiment details and the full prompt are presented in Appendix A and Appendix B, respectively.
@@ 4.1 Experimental Setup
We conducted three markets using agents based on six different API-based commercial LLMs: Claude 3.5 Sonnet, GPT-4o, Grok-2, Mistral Large 24.11, GPT-3.5, and Gemini 1.5 Pro. Each market comprised of 20 identical instances of the same LLM agent. The agents were given the same prompt and were unaware of the identities of their counter-parties. Each instance was instructed to maximize its individual profit.
For comparison, we also ran nineteen human-only markets. The average plot from these experiments are shown alongside the LLM-only Battle Royale market results in Figure 2.
4.1.1 Hypotheses
  • Rational (R): Models will not generate bubbles and trade close to the fundamental value.
  • Human (H): Models will generate bubbles similarly to humans.
  • Erratic (E): Models will perform erratically, failing to adhere to any consistent strategy, whether rational or human-like.
4.1.2 Market Analysis
@@
Claude 3.5 Sonnet 0.5361 0.0013 R 26.15
GPT-4o 0.7889 0.0497 R 22.76
Grok-2 17.3250 0.5582 H 39.43
Mistral Large 24.11 5.6944 -0.1118 H 49.96
GPT-3.5 26.3667 0.4897 H 30.44
Gemini 1.5 Pro 3.1031 -0.4009 E 45.58
MSE vs. Fundamental Value measures rationality by assessing how closely the market prices align with the fundamental value.  PCC vs. Human Market Average quantifies the similarity of LLM-driven markets to typical human market price trajectories using Pearson correlation coefficient (PCC). Portfolio Value (PV) Variance reflects strategy diversity among agents, indicating whether they adopt varied approaches or behave uniformly. The Closest Hypothesis column maps each model’s behavior to the closest hypothesized market behavior pattern.
Figure 1: Illustration of four single-model market dynamics, highlighting variations in price trajectories and agent behaviors
As shown in Table 1, Claude 3.5 Sonnet and GPT-4o achieve the lowest MSEs and smallest correlation coefficients with human price trajectories, placing them closest to the Rational (R) hypothesis. In consistently keeping prices near the fundamental value, both models behave rationally in the context of financial theory (trading only when optimal based on fundamental value). We also note that the relatively low PV Variance values for both models point to rather consistent trading profiles.
Both Grok-2 and GPT-3.5 markets demonstrated the largest deviations from fundamental as measured via MSE (least rational), with GPT-3.5 leading the way. They also recorded the two largest PCC values, suggesting a similarity to human markets. Visually (See Figure 1), Grok-2 markets show bubble formation and subsequent crash, aligning with common human market dynamics, while GPT-3.5 develops a bubble without a subsequent crash, characterized by steadily rising prices. While not the prototypical bubble, such phenomenon was also observed (albeit rarely) in human markets. We present trajectories of human markets in Figure 4 of Appendix 4. Thus, both models behave most similarly to the Human (H) hypothesis.
The Mistral Large 24.11 markets also mirror human bubble dynamics albeit with a smaller bubble size and harder crash that sees market prices dip below fundamental value in the final rounds see Figure 7 in Appendix 4. While these markets are notably more rational than humans (MSE of 5.6944 vs. 429.80 for the average human market price), the market dynamics are consistent with bubble formation and subsequent crash, aligning it most closely with the Human (H) hypothesis.
Neither rational nor human-like, Gemini 1.5 Pro exhibits a “reverse bubble,” trading well below fundamental the asset before correcting upward, reflected by its large, negative PCC value. The extent to which prices dip below fundamental was not seen in human markets nor is it rational in that agents should always be willing to purchase the asset at fundamental. Such behavior may reflect a lack of task comprehension, erratic performance, or some mixture of the two, and so this market maps most closely to the Erratic (E) hypothesis.
@@ 5 Battle Royale: Comparing LLM Market Performance
To evaluate how various LLM agents fare when competing in the same market, we conducted Battle Royale markets with a heterogeneous mixture of the models. By introducing diverse LLMs into a single environment, we aimed to explore:
  • Which LLM is most effective at maximizing profits in a competitive, multi-agent setting?
  • How a mixed-LLM market behaves, and whether it exhibits emergent phenomena when models with distinct trading strategies interact.
@@ 5.1 Experimental Setup
We conducted markets involving the same six LLM models as in the individual section. Each market included a total of 24 agents, (four agents per model). The agents operated under identical conditions and were tasked with maximizing their individual profits over 30 trading periods.
The risky asset in these markets maintained a constant fundamental value of 14 units, consistent with prior experiments.
5.2 Market Analysis
Market-Level Observations
Figure 2: Market dynamics for Battle Royale markets with average human market price for reference
Most of the Battle Royale markets (two out of three) exhibit prices near the fundamental value, indicating that these models generally converge on a “rational” pricing strategy when competing directly against one another. However, in one run, we observe a noticeable bubble that resembles the behavior typically seen in human-only markets, albeit with a lower peak price. Notably, there is a clear uptick in trading volume near the end of the bubble run, which mirrors human experiments where participants rush to adjust their positions before the market’s final redemption. Overall, these patterns align most closely with Rational Hypothesis (R), where agents trade at or near fundamental value. Nevertheless, the results also demonstrate that even ostensibly rational agents can occasionally generate bubble-like dynamics under competitive, multi-agent conditions.
@@ 5.2 Market Analysis
Model Avg Portfolio Value PV Variance
Mistral Large 24.11 689.56 16.91
Gemini 1.5 Pro 688.49 20.05
Claude 3.5 Sonnet 680.36 25.83
GPT-3.5 674.42 12.08
GPT-4o 671.36 17.30
Grok-2 668.51 27.01

### v3

# LLM Agents Do Not Replicate Human Market Traders: Evidence from Experimental Finance

@@ Abstract
In this study, we compare Large Language Models (LLMs) with human traders in a classic experimental-finance paradigm where prices are determined endogenously. Using a well-established asset-trading design, we run homogeneous (mono-agent) markets with single-model LLM agents and heterogeneous “Battle Royale” markets with multiple LLM models. Our findings reveal that LLMs generally exhibit a “textbook-rational” approach, pricing the asset near its fundamental value and showing only a muted tendency toward bubble formation, whereas humans deviate substantially and consistently generate bubbles. Additional treatments, including dividend shocks and repeated-exposure (‘experienced’) runs, show that these differences persist across various experimental settings. Further analyses of LLM-generated strategy text indicate lower variance, reduced bias, and stronger reliance on fundamentals relative to humans’ more heuristic-driven trading. These results highlight the risk of using LLM-agents to model human-driven market phenomena, as key behavioral features such as large, emergent bubbles are not reproduced.
@@ 1 Introduction
Financial markets are complex, dynamic coordination systems shaped by individual decisions made under uncertainty. While traditional financial theory assumes rational behavior, decades of behavioral finance research has revealed persistent biases, such as herding and overconfidence, that can give rise to mispricings and asset bubbles. These deviations underscore the challenges of achieving optimal outcomes in uncertain and evolving market environments.
The rise of agentic systems, particularly those driven by Large Language Models (LLMs), introduces a novel and previously unexplored dimension to this problem. As LLMs increasingly participate in domains traditionally navigated by humans, including financial systems and markets, understanding their behavioral tendencies has become critical. Questions such as; Do LLM-agents exhibit similar behavioral biases, behave more rationally, or fall victim to entirely different biases not seen in humans?
This paper presents the first comprehensive study comparing ”out-of-the-box” LLM and human behavior within a single experimental finance paradigm. We adopt a classical trading environment where participants trade a risky asset with a controlled fundamental value and examine the behavior of LLM-agents in both homogeneous (mono-agent) markets, populated entirely by instances of the same model, and heterogeneous ”Battle Royale” markets, where agents based on different base LLMs compete. These experimental markets are entirely endogenous, with prices determined solely by the buy-and-sell orders submitted by participants, and without any external price anchor or market maker. This enables us to dive deeper into how the underlying agent behavior affects the market, as we can control all market parameters.
Previous research offers conflicting perspectives on AI-human behavioral alignment. Some works suggest a significant difference between human and LLM interactions (Anlló et al. (2024)), while others indicate that LLM behavior could supplement or even substitute for human studies (Leng (2024); Horton (2023); Jia et al. (2024)). We address this tension in the literature through rigorous experimental comparison in experimental finance, a domain that has long served as a testbed for understanding financial decision-making under risk and uncertainty.
@@
Our findings reveal clear behavioral differences between LLM and human-driven markets. LLM markets demonstrate substantially more “textbook-rational” behavior than their human counterparts. Human markets consistently generate significant price bubbles and deviations from the asset’s fundamental value, while LLM-agents generally trade much closer to fundamental value and exhibit markedly reduced trading variance. We also find that LLMs appear to incorporate less bias into their future price forecasts and employ strategies more focused on fundamental value, rather than strategies commonly used by humans. These results challenge the notion that ”out-of-the-box” LLMs can faithfully replicate human market dynamics, particularly in terms of the emergence of market-level phenomena such as bubbles and crashes. As agentic systems increasingly participate in and influence real financial markets, understanding these behavioral divergences becomes essential for effective regulation, risk management, and market design.
@@ 2 Background
Numerous works have studied LLMs in strategic environments such as social dilemmas and iterated games Brookins & DeBacker (2024); Horton (2023); Akata et al. (2023). In finance, most existing studies focus on empirical applications—LLMs trading in real markets or analyzing financial data (e.g., Xiao et al. (2024); Li et al. (2024); Yu et al. (2024a); Yun (2024); Liu & Lo (2025); Ding et al. (2024); Abdelsamie & Wang (2024); Chen (2025); Zhou et al. (2025)). However, to our knowledge, no prior work has analyzed LLM behavior in experimental markets where prices are endogenously generated by agent interaction (see Appendix C for a more comprehensive review).
In finance, LLMs have regularly been applied to real-world scenarios. Papers such as; Zhang et al. (2023); Bond (2023); Kirtac & Germano (2024) explored LLMs’ ability to process financial sentiment and predict market movements. Fatouros et al. (2024) introduced MarketSenseAI, which leverages GPT-4 for stock selection, illustrating the utility of LLMs in navigating complex financial data. While these studies demonstrate the promise of LLMs in financial settings, there is still a need to understand how these agents perform in experimental financial settings where strategic interactions drive outcomes.
The study of market dynamics in controlled experimental environments has revealed critical insights into behavioral biases and inefficiencies. For instance, Smith et al. (1988) demonstrated the formation of bubbles and crashes in experimental markets, even under conditions of perfect information. Subsequent work by Bostian & Holt (2009) highlighted how individual behaviors, such as overconfidence and herding, contribute to these phenomena. Incorporating AI agents into such paradigms presents a unique opportunity to evaluate whether these systems replicate, mitigate, or exacerbate market inefficiencies.
@@ 3 Experiment
We focus on a classic experimental finance paradigm with a constant fundamental value, the mathematical price of an asset; see Equation B in the Appendix. We adopt the experimental design from Smith et al. (2014) and Holt et al. (2017), which involves an endogenous experimental asset market where subjects trade a risky asset (STOCK) in an open-call market for 30 periods operated using the oTree experiment platform (Chen et al. (2016)). Participants are given 4 shares of STOCK and 100 units of CASH. Participants have two options: either invest CASH in STOCK, which in each trading period pays out dividends of either 0.4 or 1 units of CASH with equal probability, or hold CASH, which generates interest at 5% each trading period. The price is determined solely by the trading activity of the experimental market participants.
After the experiment, the risky asset is redeemed for a fixed value of 14 units of CASH, representing the fundamental value of the STOCK (See Appendix B for full formulation).
Because of these parameter decisions (i.e., dividend schedule, CASH interest rate, and redemption price), the fundamental value remains constant throughout the entirety of the experiment, enabling us to objectively classify when the market is in a “bubble” state Bostian et al. (2005). This differs starkly from applied financial studies, where the fundamental value of a real-world asset is unknown. Further discussion of the experimental details can be found in Appendices A.
This task was assigned to both LLM agents and human traders in separate markets. In adapting the task to LLM-agents, our primary goal was to best emulate the experiment presented to human subjects. The description of the task setting mirrors the original instruction video, shown to human participants (See Appendix H.2). In every round of the experiment, each agent simultaneously submits orders to the market as well as expectations regarding current and future market prices.
A common challenge in adapting multi-round games to LLM-agents is the persistence of memory between rounds, which is critical to learning and adapting strategies. Building off previous works (Fish et al. (2024)), we allow agents to communicate with future versions of themselves by allowing them to read and write their own “Insights” and “Thoughts” at each round. This structure incorporates Chain-of-Thought (CoT) reasoning (Wei et al. (2022)), which has been shown to help elicit the reasoning ability of LLMs, providing an interpretable window into how agents perceive and interact with the market. Further prompting details are presented in Appendix I.
To frame our analysis, we categorized potential outcomes into three groups: rational (R), where models do not generate bubbles and trade close to the fundamental value; human (H), where models generate bubbles in a manner similar to humans; and erratic (E), where models perform inconsistently, failing to adhere to any consistent strategy, whether rational or human-like.
@@ 4.1 Experimental Setup
We conducted three markets using agents based on six different API-based commercial LLMs: Anthropic’s Claude-3.5-Sonnet Anthropic (2024), OpenAI’s GPT-3.5 and GPT-4o OpenAI (2025), xAI’s Grok-2 xAI (2024), Mistral-Large AI’s Mistral-Large AI (2024), and Google’s Gemini-1.5 Pro DeepMind (2024). Each market consisted of 20 identical instances of the same LLM agent. The agents were given the same prompt and were unaware of the identities of their counterparts. Each instance was instructed to maximize its profit.
For comparison, we also ran nineteen human-only markets, with identical experimental conditions (though the humans also conducted a short risk-elicitation task in between each round). Participants were a mix of [UNIVERSITY] students and staff, as well as online participants recruited through prolific.com. The human participants were given instructions that mirror the LLM prompts; full instructions are shown in Appendix section H.2. The average plot from these experiments is shown alongside the LLM-only Battle Royale market results in Figure 5.1.
4.1.1 Single-Model Market Analysis
As shown in Table 1, Claude-3.5 Sonnet and GPT-4o yield the lowest MSEs and nearly zero correlation with human price paths, keeping prices close to fundamentals and mapping to the Rational (R) hypothesis. Both also show low portfolio variance, suggesting consistent strategies across instances.
By contrast, Grok-2 and GPT-3.5 show the largest deviations from fundamentals and the highest correlations with human markets. Grok-2 exhibits bubble–crash cycles, while GPT-3.5 generates a monotonic run-up without correction (Figure 1), a pattern occasionally seen in humans.
Mistral-Large also produces bubbles but with smaller peaks and sharper crashes, dipping below fundamentals late in the session. While more rational in absolute terms (MSE $\approx$ 5.7 vs. 429.8 for humans), its dynamics align with the Human (H) hypothesis.
Finally, Gemini-1.5 Pro departs from both rational and human-like behavior, trading well below fundamentals before correcting upward. This so-called “reverse bubble” (negative PCC) suggests erratic dynamics rather than coherent strategy, mapping most closely to the Erratic (E) hypothesis.
@@ 4.1.1 Single-Model Market Analysis
Claude-3.5-Sonnet 0.536 0.001 R 26.15
GPT-4o 0.789 0.050 R 22.76
Grok-2 17.325 0.558 H 39.43
Mistral-Large 5.694 -0.112 H 49.96
GPT-3.5 26.367 0.490 H 30.44
Gemini-1.5 Pro 3.103 -0.401 E 45.58
Notes: MSE vs. Fundamental Value measures rationality by assessing how closely the market prices align with the fundamental value. PCC vs. Human Market Average quantifies the similarity of LLM-driven markets to typical human market price trajectories using the Pearson correlation coefficient (PCC). Portfolio Value (PV) Variance reflects strategy diversity among agents, indicating whether they adopt varied approaches or behave uniformly. The Closest Hypothesis column maps each model’s behavior to the closest hypothesized market behavior pattern, based on a mix of quantitative and qualitative comparisons. R if MSE around fundamental was below 1, and H if we observed a bubble formation and crash, E if the market behavior was nonsensical (reverse bubble).
Figure 1: Mono-Agent Price Charts compared to Humans
None of the LLM-agents generate market bubbles similar in magnitude to humans.
@@ 5 Battle Royale: Comparing LLM Market Performance
To evaluate how various LLM agents fare when competing in the same market, we conducted Battle Royale market experiment with a heterogeneous mixture of the models. By introducing diverse LLMs into a single environment, we aimed to explore:
  • Which LLM is most effective at maximizing profits in a competitive, multi-agent setting? Which Agent is the best trader?
  • How a mixed-LLM market behaves, and whether it will exhibit emergent phenomena when models with distinct trading strategies interact?
  • Whether heterogeneous LLM interpretation and strategy differences are enough to generate financial bubbles?
@@ 5.1 Experimental Setup
We conducted markets involving the same six LLM models as in the Single-Model section. Each market included a total of 24 agents (four agents per model). The agents operated under identical conditions and were tasked with maximizing their profits over 30 trading periods. The risky asset in these markets maintained a constant fundamental value of 14 units, consistent with prior experiments. Figure 2 shows the three ’Battle Royale’ market results.
Figure 2: Battle Royale Market Performance with Human Comparison
@@
Model Avg Portfolio Value $\pm$ 1 STD
Mistral-Large 689.56 $\pm$ 4.11
Gemini-1.5 Pro 688.49 $\pm$ 4.48
Claude-3.5 Sonnet 680.36 $\pm$ 5.08
GPT-3.5 674.42 $\pm$ 3.48
GPT-4o 671.36 $\pm$ 17.30
Grok-2 668.51 $\pm$ 27.01